# Endomorphism algebras of abelian varieties with large cyclic 2-torsion field over a given field

### v3

@@ Abstract.
In this article we study the endomorphism algebras of abelian varieties $A$ defined over a given number field $K$ with large cyclic 2-torsion fields. A key step in doing so is to provide criteria for all the endomorphisms of $A$ to be defined over $K(A[2])$, the field generated by its 2-torsion.
When $K=\mathbb{Q}$ and $\operatorname{{\rm Gal}}(\mathbb{Q}(A[2])/\mathbb{Q})$ is cyclic of prime order $p=2\dim(A)+1$, we prove the somewhat surprising result that there are only finitely many possibilities for the geometric endomorphism algebra $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$. In fact, when $\dim(A)\not\in\{3,5,9,21,33,81\}$, we show $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is a proper subfield of the $p$-th cyclotomic field.
In particular, when $g=2$, $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is isomorphic to either $\mathbb{Q}$ or $\mathbb{Q}(\sqrt{5})$.
@@ 1. Introduction
Faltings’ Isogeny Theorem tells us that for an abelian variety $A$, defined over a number field $K$, we can recover its ring of endomorphisms defined over $K$, $\operatorname{{\rm End}}_{K}(A)$, tensored by $\mathbb{Z}_{\ell}$ (where $\ell$ is a prime) from looking at the action of the absolute Galois group $G_{K}\coloneqq\operatorname{{\rm Gal}}(\bar{K}/K)$ on its $\ell$-adic Tate module $T_{\ell}(A)$. More precisely, Faltings’ Isogeny Theorem tells us that the natural injection
@@
is not an isomorphism. However, when the action of $G_{K}$ on $A[\ell]$ is particularly ‘large’ one might hope that $\operatorname{{\rm End}}_{K}(A[\ell])\;(=\operatorname{{\rm End}}_{\bar{K}}(A[\ell])^{G_{K}})$ is small enough to place significant restrictions on $\operatorname{{\rm End}}_{K}(A)\otimes\mathbb{Z}/\ell\mathbb{Z}$ and hence $\operatorname{{\rm End}}_{K}(A)$.
@@
In the cases considered by Zarhin, all of the Galois groups are insoluble and act 2-transitively on the roots of $f$. In [Goo21] the author considered restrictions on $\operatorname{{\rm End}}_{K}(A)$ and $\operatorname{{\rm End}}(A)\coloneqq\operatorname{{\rm End}}_{\bar{K}}(A)$ when $\operatorname{{\rm Gal}}(K(A[\ell])/K)$ contained merely an element of large prime order, with no restrictions on $K$.
However, considering the case of elliptic curves, for example, where the possible endomorphism algebras depend solely on $[K:\mathbb{Q}]$, it is natural to impose conditions on $K$. The main advantage in doing so is that this allows for more arithmetic input, whereas the proofs of the above results are heavily representation theoretic.
@@
As an example, we obtain the somewhat surprising [^1] At least in the author’s opinion. result (where $\zeta_{n}$ denotes a primitive $n$-th root of unity):
@@
  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\bar{K}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
@@
Intuitively, the reason for this discrepancy is already exemplified in the above example: $\operatorname{{\rm Gal}}(K(A[2])/K)$ being small should force the endomorphisms to be defined over a small extension of $K$. To make this rigorous, we need to study the minimal extension $L/K$ over which $A$ obtains all its endomorphisms. This field was christened in [GK17] as the endomorphism field of $A$
An important result in understanding this field was obtained by Silverberg, who showed it is contained in $K(A[n])$ for $n\geq 3$ [Sil92, Theorem 2.4]. Bounds on $[L\colon K]$ were then studied in [GK17] and also in [FKRS12, FKS21] via its connection to Sato-Tate groups.
@@
Theorem 1.2 ($=$ Theorem 2.1).
Suppose $E=\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is a (finite) Galois extension of $\mathbb{Q}$ and $L\nsubseteq K(A[2])$. The following hold:
  • $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ has a non-trivial normal elementary abelian 2-subgroup;
  • if $\operatorname{{\rm End}}(A)$ is 2-maximal in $E$, then $2$ is wildly ramified in $E/\mathbb{Q}$.
In particular, if $E/\mathbb{Q}$ is Galois, $\operatorname{{\rm End}}(A)$ is a 2-maximal order and $2$ is not wildly ramified, then $L\subseteq K(A[2])$.
Recall that an order $\mathcal{O}\subseteq\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is said to be 2-maximal if for any other order $\mathcal{O}^{\prime}\subseteq\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ satisfying $\mathcal{O}\otimes\mathbb{Z}_{2}\subseteq\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}$ we have $\mathcal{O}\otimes\mathbb{Z}_{2}=\mathcal{O}^{\prime}\otimes\mathbb{Z}_{2}$.
Example 2.2 shows that the condition $\operatorname{{\rm End}}(A)$ is 2-maximal cannot be removed.
Finally, we would like to end the introduction with a challenge [^2] See also Question 2.12 for the reader which we hope will provide some impetus for studying moduli spaces of abelian surfaces. To do so let us state the following corollary of Theorems 1.1,1.2:
Corollary 1.3 ($=$ Corollary 2.8).
@@
Challenge 1.4.
Give an independent proof of Corollary 1.3 using moduli spaces of abelian surfaces.
@@ Acknowledgements.
The author thanks Tim Dokchitser, Francesc Fité, Xavier Guitart, Jef Laga, Jeremy Rickard, Víctor Rotger, Ciaran Schembri, John Voight and David Zywina for useful comments, questions and discussions.
The author thanks the University of Bristol and the Université Clermont Auvergne for their hospitality during the period this paper was written. Where the author was funded by the EPSRC (grant number EP/N509619/1) and an early career fellowship from the LMS respectively. Finally, the author thanks Universitat de Barcelona where final revisions to this paper were made.
@@
Throughout $A$ will denote an abelian variety of dimension $g$ defined over a number field $K$ (often $\mathbb{Q}$). Moreover, we denote by $L$ the smallest extension of $K$ over which all the endomorphisms of $A$ are defined.
We write $\operatorname{{\rm End}}_{F}(A)$ for the ring of endomorphisms of $A$ defined over $F/K$, $\operatorname{{\rm End}}(A)\coloneqq\operatorname{{\rm End}}_{\bar{K}}(A)$ and $\operatorname{{\rm End}}_{F}^{0}(A)\coloneqq\operatorname{{\rm End}}_{F}(A)\otimes\mathbb{Q}$.
@@ 2. Proof of the main theorems
Silverberg has shown the endomorphism field $L/K$ of an abelian variety $A/K$ is contained in the $n$-torsion field for every $n\geq 3$ [Sil92, Theorem 2.4]. The case $n=4$ implies $\operatorname{{\rm Gal}}(L/L\cap K(A[2]))$ is an elementary abelian 2-group, since $\operatorname{{\rm Gal}}(K(A[4])/K(A[2]))$ is too, for details see [Goo21, Proposition 3.9]. Below we provide conditions on $\operatorname{{\rm End}}(A)$ which guarantee $L\subseteq K(A[2])$.
@@
Suppose $E=\operatorname{{\rm End}}^{0}(A)$ is a (finite) Galois extension of $\mathbb{Q}$ and $L\nsubseteq K(A[2])$. The following hold:
  • $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ has a non-trivial normal elementary abelian 2-subgroup;
  • if $\operatorname{{\rm End}}(A)$ is 2-maximal in $E$, then $2$ is wildly ramified in $E/\mathbb{Q}$.
In particular, if $E/\mathbb{Q}$ is Galois, $\operatorname{{\rm End}}(A)$ is a 2-maximal order and $2$ is not wildly ramified, then $L\subseteq K(A[2])$.
@@
If $\operatorname{{\rm Gal}}(L/K)$ acts faithfully on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$, then we may recuperate Lemma 2.1 in [Sil92] with $N=2$, $\Lambda=\operatorname{{\rm End}}(A)$ and $G$ the image of $\operatorname{{\rm Gal}}(L/K)$ in $\operatorname{{\rm Aut}}(\Lambda)$. Following through the proofs of Propositions 2.2, 2.3 and Theorem 2.4 in [Sil92] shows $L\subseteq K(A[2])$. Hence we may suppose the action of $\operatorname{{\rm Gal}}(L/K)$ on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ has a non-trivial element $\gamma$ in its kernel.
Since $\operatorname{{\rm Gal}}(L/K)$ acts on $\operatorname{{\rm End}}(A)$ by field automorphisms of $\operatorname{{\rm End}}^{0}(A)$, we may view its image as a subgroup of $\operatorname{{\rm Gal}}(E/\mathbb{Q})$. Identify $\gamma$ with its image in $\operatorname{{\rm Gal}}(E/\mathbb{Q})$. Let $d$ be the rank of $\operatorname{{\rm End}}(A)$ as a free $\mathbb{Z}$-module and view $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ as a (finite) subgroup of $\operatorname{{\rm GL}}_{d}(\mathbb{Z})$ via its action on $\operatorname{{\rm End}}(A)$. The reduction map $\operatorname{{\rm GL}}_{d}(\mathbb{Z})\rightarrow\operatorname{{\rm GL}}_{d}(\mathbb{Z}/4\mathbb{Z})$ is injective on finite groups and the kernel of $\operatorname{{\rm GL}}_{d}(\mathbb{Z}/4\mathbb{Z})\rightarrow\operatorname{{\rm GL}}_{d}(\mathbb{Z}/2\mathbb{Z})$ is an elementary abelian $2$-group (for more details see [Goo21, Prop. 3.9]). Thus the kernel of the map $\operatorname{{\rm Gal}}(E/\mathbb{Q})\rightarrow\operatorname{{\rm Aut}}(\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z})\cong\operatorname{{\rm GL}}_{d}(\mathbb{Z}/2\mathbb{Z})$ is an elementary abelian $2$-group which is non-trivial since it contains $\gamma$.
Assume now $\operatorname{{\rm End}}(A)$ is a 2-maximal order in $E$. As $\gamma$ acts trivially on the order $\operatorname{{\rm End}}(A)$ modulo $2$, it also acts trivially on $\operatorname{{\rm End}}(A)$ modulo every prime above $2$ in $E$. Since $\operatorname{{\rm End}}(A)$ is 2-maximal, we find $\gamma$ belongs to the inertia group $I_{\mathfrak{p}}$ for every prime $\mathfrak{p}$ above 2 in $E$. By the above $\gamma$ has order two, and hence $2$ is wildly ramified in $E/\mathbb{Q}$. ∎
@@
Let $A/K$ be an abelian variety of dimension $g\geq 1$ with $p=2g+1$ prime. Suppose $\mathfrak{q}$ is a prime of bad reduction for $A$ and there is an element of order $p$ in the image of the inertia group $I_{\mathfrak{q}}$ on $A[\ell]$ for some $\mathfrak{q}\nmid\ell$. Then either
  • $p$ does not divide $[L:K]$ and $\operatorname{{\rm End}}^{0}(A)$ is a subfield of $\mathbb{Q}(\zeta_{p})$; or
  • $p$ divides $[L:K]$, $g\geq 3$ and $A$ is isogenous over $\bar{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield of $\mathbb{Q}(\zeta_{p})$.
@@
Suppose $p$ divides $[L:\mathbb{Q}]$. Then by Theorem 2.5 of [Goo21], $g\geq 3$ and $A$ is isogenous over $\bar{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield $F$ of $\mathbb{Q}(\zeta_{p})$. Hence we may suppose $p$ does not divide $[L:K]$.
The criterion of Néron-Ogg-Shafarevich implies that the image of the inertia group $I_{\mathfrak{q}}$ on the Tate module $T_{\ell}(A)$ for any prime $\ell$, not divisible by $\mathfrak{q}$, contains an element $\tau$ of order $p$. Furthermore, as the trace of $\tau$ is an integer [ST68, Thm. 2] its eigenvalues are the primitive $p$-th roots of unity (owing to the fact $\tau$ has at most $12g=p-$ 1 distinct eigenvalues).
By Dirichlet’s theorem on arithmetic progressions, we can find a prime $\ell$, not divisible by $\mathfrak{q}$, which is a primitive root modulo $p$. The reduction of $\tau$ modulo $\ell$ lands in $\operatorname{{\rm Gal}}(K(A[\ell])/K)$ and has order $p$. This allows us to apply [Goo21, Theorem 2.9] and deduce $E\coloneqq\operatorname{{\rm End}}^{0}(A)$ is a field.
As $p$ does not divide $[L:K]$, the element $\tau$ lies in the image of $G_{L}$. The decomposition $E\otimes\mathbb{Q}_{\ell}=\prod_{\lambda|\ell}E_{\lambda}$ induces a decomposition $V_{\ell}(A)=\prod_{\lambda|\ell}V_{\lambda}(A)$ giving representations $G_{L}\rightarrow\operatorname{{\rm GL}}_{n}(E_{\lambda})$ where $n=\frac{2g}{[E:\mathbb{Q}]}$, see §2 of [Rib76] for further details. Let $\tau_{\lambda}\in\operatorname{{\rm GL}}_{n}(E_{\lambda})$ be the projection of $\tau$ onto $V_{\lambda}(A)$. By the above, the eigenvalues of $\tau_{\lambda}$ are distinct primitive $p$-th roots of unity. Taking the trace of $\tau_{\lambda}$ we deduce $E_{\lambda}$ contains a subfield of $\mathbb{Q}(\zeta_{p})$ of degree $[E:\mathbb{Q}]$.
As $\mathbb{Q}(\zeta_{p})$ has a unique subfield $F$ of degree $[E:\mathbb{Q}]$ and for all, but possibly finitely many $\lambda$, $E_{\lambda}$ contains $F$, the Chebotarev Density Theorem implies $E$ contains $F$. But $[E:\mathbb{Q}]=[F:\mathbb{Q}]$, thus in other words, $E\subseteq\mathbb{Q}(\zeta_{p})$. ∎
@@
  • $p$ does not divide $[L:K]$ and $\operatorname{{\rm End}}^{0}(A)$ is a subfield of $\mathbb{Q}(\zeta_{p})$; or
  • $p$ divides $[L:K]$, $g\geq 3$ and $A$ is isogenous over $\bar{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield of $\mathbb{Q}(\zeta_{p})$.

### v4

@@ Abstract.
In this article we study the endomorphism algebras of abelian varieties $A$ defined over a given number field $K$ with large cyclic 2-torsion fields. A key step in doing so is to provide criteria for all the endomorphisms of $A$ to be defined over $K(A[2])$, the field extension generated by its 2-torsion.
When $K=\mathbb{Q}$ and $\operatorname{{\rm Gal}}(\mathbb{Q}(A[2])/\mathbb{Q})$ is cyclic of prime order $p=2\dim(A)+1$, we prove that there are only finitely many possibilities for the geometric endomorphism algebra $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$. In fact, when $\dim(A)\not\in\{3,5,9,21,33,81\}$, we show $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is a proper subfield of the $p$-th cyclotomic field. In particular, when $g=2$, $\operatorname{{\rm End}}(A)\otimes\mathbb{Q}$ is isomorphic to either $\mathbb{Q}$ or $\mathbb{Q}(\sqrt{5})$.
@@ 1. Introduction
Faltings’ Isogeny Theorem [Fal83, p. 360, Satz 4] tells us that for an abelian variety $A$, defined over a number field $K$, we can recover its ring of endomorphisms defined over $K$, $\operatorname{{\rm End}}_{K}(A)$, tensored by $\mathbb{Z}_{\ell}$ (where $\ell$ is a prime) from looking at the action of the absolute Galois group $G_{K}\coloneqq\operatorname{{\rm Gal}}(\overline{K}/K)$ on its $\ell$-adic Tate module $T_{\ell}(A)$. More precisely, Faltings’ Isogeny Theorem tells us that the natural injection
@@
is not an isomorphism. However, when the action of $G_{K}$ on $A[\ell]$ is particularly ‘large’ one might hope that $\operatorname{{\rm End}}_{K}(A[\ell])\;(=\operatorname{{\rm End}}_{\overline{K}}(A[\ell])^{G_{K}})$ is small enough to place significant restrictions on $\operatorname{{\rm End}}_{K}(A)\otimes\mathbb{Z}/\ell\mathbb{Z}$ and hence $\operatorname{{\rm End}}_{K}(A)$.
@@
In the cases considered by Zarhin, all of the Galois groups are insoluble and act 2-transitively on the roots of $f$. In [Goo21] the author considered restrictions on $\operatorname{{\rm End}}_{K}(A)$ and $\operatorname{{\rm End}}(A)\coloneqq\operatorname{{\rm End}}_{\overline{K}}(A)$ when $\operatorname{{\rm Gal}}(K(A[\ell])/K)$ contained merely an element of large prime order, with no restrictions on $K$.
However, considering the case of elliptic curves, for example, where the possible endomorphism algebras depend solely on $K\colon\mathbb{Q}$], it is natural to impose conditions on $K$. The main advantage in doing so is that this allows for more arithmetic input, whereas the proofs of the above results are heavily representation theoretic.
@@
As an example, we obtain (where $\zeta_{n}$ denotes a primitive $n$-th root of unity):
@@
  • $p\in\{7,11,19,43,67,163\}$ and $A$ is isogenous over $\overline{K}$ to the power of an elliptic curve with complex multiplication by $\mathbb{Q}(\sqrt{-p})$.
@@
Intuitively, the reason for this discrepancy is already exemplified in the above example: $\operatorname{{\rm Gal}}(K(A[2])/K)$ being small should force the endomorphisms to be defined over a small extension of $K$. To make this rigorous, we need to study the minimal extension $L/K$ over which $A$ obtains all its endomorphisms. This field was christened in [GK17] as the endomorphism field of $A$ .
An important result in understanding this field was obtained by Silverberg, who showed it is contained in $K(A[n])$ for $n\geq 3$ [Sil92, Thm. 2.4]. Bounds on $L\colon K]$ were then studied in [GK17] and also in [FKRS12, FKS25] via its connection to Sato-Tate groups.
@@
Theorem 1.2 (== Theorem 2.1).
Suppose $E\cong\operatorname{{\rm End}}^{0}(A)$ is a (finite) Galois extension of $\mathbb{Q}$ and $L\nsubseteq K(A[2])$. Then $E\colon\mathbb{Q}$] is even.
If moreover $\operatorname{{\rm End}}(A)$ is 2-maximal in $E$, then $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ has a non-trivial normal elementary abelian 2-subgroup which is contained in the intersection of inertia groups $\cap_{\mathfrak{p}|2}I_{\mathfrak{p}}$ where $\mathfrak{p}$ runs over all primes above $2$ in $E$.
In particular, if $E/\mathbb{Q}$ is Galois, $\operatorname{{\rm End}}(A)$ is a 2-maximal order and $2$ is not wildly ramified in $E$, then $L\subseteq K(A[2])$.
Recall that an order $\mathcal{O}$ in a number field $E$ is said to be 2-maximal if any of the equivalent conditions hold: its index in the maximal order $\mathcal{O}_{E}$ is odd; $\mathcal{O}\otimes\mathbb{Z}_{2}=\mathcal{O}_{E}\otimes\mathbb{Z}_{2}$; or $\mathcal{O}\otimes\mathbb{Z}_{(2)}=\mathcal{O}_{E}\otimes\mathbb{Z}_{(2)}$ [Voi21, Lem. 9.5.3, Lem. 9.6.7]. Example 2.2 shows that the condition $\operatorname{{\rm End}}(A)$ is 2-maximal cannot be removed.
Finally, let us underline the following corollary of Theorems 1.1,1.2.
Corollary 1.3 (== Corollary 2.8).
@@ Acknowledgements.
The author thanks Tim Dokchitser, Francesc Fité, Xavier Guitart, Jef Laga, Jeremy Rickard, Víctor Rotger, Ciaran Schembri, John Voight, David Zywina and the anonymous referee for useful comments, questions and discussions.
The author thanks the University of Bristol and the Université Clermont Auvergne for their hospitality during the period this paper was written. Where the author was funded by the EPSRC (grant number EP/N509619/1) and an early career fellowship from the LMS respectively. Finally, the author thanks Universitat de Barcelona where final revisions to this paper were made whilst he was supported by the Spanish Ministry of Science and Innovation via the grant “Abelian varieties, L-functions, and rational points” (code PID2022-137605NB-I00).
@@
Throughout $A$ will denote an abelian variety of dimension $g$ defined over a number field $K$. Moreover, fixing an algebraic closure of $\overline{K}$ of $K$, we denote by $L$ the smallest extension of $K$ contained in $\overline{K}$ over which all the endomorphisms of $A$ are defined.
For an extension $F/K$ contained in $\overline{K}$, we write $\operatorname{{\rm End}}_{F}(A)$ for the ring of endomorphisms of $A$ defined over $F$, $\operatorname{{\rm End}}(A)\coloneqq\operatorname{{\rm End}}_{\overline{K}}(A)$ and $\operatorname{{\rm End}}_{F}^{0}(A)\coloneqq\operatorname{{\rm End}}_{F}(A)\otimes\mathbb{Q}$. Letting $G_{F}$ denote the absolute Galois group of $F$, we have $\operatorname{{\rm End}}(A)^{G_{F}}=\operatorname{{\rm End}}_{F}(A)$. We write $\mathrm{Cl}(F)$ for the class group of $F$ and $\mathrm{Cl}(F)$ for its order.
@@ 2. Proof of the main theorems
Silverberg has shown the endomorphism field $L$ of an abelian variety $A/K$ is contained in the $n$-torsion field for every $n\geq 3$ [Sil92, Thm. 2.4]. The case $n=4$ implies $\operatorname{{\rm Gal}}(L/L\cap K(A[2]))$ is an elementary abelian 2-group, since $\operatorname{{\rm Gal}}(K(A[4])/K(A[2]))$ is too, for details see [Goo21, Prop. 3.9]. Below we provide conditions on $\operatorname{{\rm End}}(A)$ which guarantee $L\subseteq K(A[2])$.
@@
Suppose $E\cong\operatorname{{\rm End}}^{0}(A)$ is a (finite) Galois extension of $\mathbb{Q}$ and $L\nsubseteq K(A[2])$. Then $E\colon\mathbb{Q}$] is even.
If moreover $\operatorname{{\rm End}}(A)$ is 2-maximal in $E$, then $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ has a non-trivial normal elementary abelian 2-subgroup which is contained in the intersection of inertia groups $\cap_{\mathfrak{p}|2}I_{\mathfrak{p}}$ where $\mathfrak{p}$ runs over all primes above $2$ in $E$.
In particular, if $E/\mathbb{Q}$ is Galois, $\operatorname{{\rm End}}(A)$ is a 2-maximal order and $2$ is not wildly ramified in $E$, then $L\subseteq K(A[2])$.
@@
Suppose $\operatorname{{\rm Gal}}(L/K)$ acts faithfully on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$. Recall that the natural map $\operatorname{{\rm End}}(A)\rightarrow\operatorname{{\rm End}}(A[2])$ induces an injection $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}\hookrightarrow\operatorname{{\rm End}}(A[2])$. The group $\operatorname{{\rm Gal}}(\overline{K}/K(A[2]))$ acts trivially on $\operatorname{{\rm End}}(A[2])$ and hence trivially on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ as well. Since $\operatorname{{\rm Gal}}(L/K)$ acts faithfully on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$, we find $L\subseteq K(A[2])$. As we are assuming $L\not\subseteq K(A[2])$, we deduce that the action of $\operatorname{{\rm Gal}}(L/K)$ on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}$ has a non-trivial element $\gamma$ in its kernel.
The action of $\operatorname{{\rm Gal}}(L/K)$ on $\operatorname{{\rm End}}(A)$ is faithful, allowing us to view $\operatorname{{\rm Gal}}(L/K)$ as a subgroup of $\operatorname{{\rm Aut}}_{\mathbb{Z}}(\operatorname{{\rm End}}(A))\cong\operatorname{{\rm GL}}_{d}(\mathbb{Z})$ where $d$ is the rank of $\operatorname{{\rm End}}(A)$ viewed as a free $\mathbb{Z}$-module. The action also extends to field automorphisms of $\operatorname{{\rm End}}^{0}(A)\cong E$, thus we may also view $\operatorname{{\rm Gal}}(L/K)$ as a subgroup of $\operatorname{{\rm Gal}}(E/\mathbb{Q})$. Any finite subgroup in the kernel of the reduction map $\operatorname{{\rm GL}}_{d}(\mathbb{Z})\rightarrow\operatorname{{\rm GL}}_{d}(\mathbb{Z}/2\mathbb{Z})$ is an elementary abelian 2-group [GL06, p. 151, Lem. 9]. As $\operatorname{{\rm Aut}}_{\mathbb{Z}/2\mathbb{Z}}(\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z})\cong\operatorname{{\rm GL}}_{d}(\mathbb{Z}/2\mathbb{Z})$, we deduce that $\{\sigma\in\operatorname{{\rm Gal}}(L/K)|(\sigma-1)\operatorname{{\rm End}}(A)\subseteq 2\cdot\operatorname{{\rm End}}(A)\}$ is an elementary abelian $2$-group which is non-trivial since it contains $\gamma$. Hence $E\colon\mathbb{Q}$] is divisible by 2.
Assume now $\operatorname{{\rm End}}(A)$ is a 2-maximal order in $E$. Hence $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{2}\cong\mathcal{O}_{E}\otimes\mathbb{Z}_{2}$ where $\mathcal{O}_{E}$ is the maximal order of $E$. As $\operatorname{{\rm Gal}}(E/\mathbb{Q})$ acts on $\mathcal{O}_{E}$, it also acts on $\mathcal{O}_{E}\otimes\mathbb{Z}_{2}$ and thus also on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{2}$ via the above isomorphism. The kernel $H$ of the composite map $\operatorname{{\rm Gal}}(E/\mathbb{Q})\hookrightarrow\operatorname{{\rm Aut}}_{\mathbb{Z}_{2}}(\operatorname{{\rm End}}(A)\otimes\mathbb{Z}_{2})\rightarrow\operatorname{{\rm Aut}}_{\mathbb{Z}/2\mathbb{Z}}(\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z})$ is an elementary abelian 2 group [GL06, p. 151, Lem. 9] [^1] The statement they give does not apply directly to $\mathbb{Z}_{2}$, however the proof carries over verbatim., which is non-trivial since it contains $\gamma$.
Write $2\mathcal{O}_{E}=\prod_{i}\mathfrak{p}^{e_{i}}_{i}$ where each $\mathfrak{p}_{i}$ is a prime ideal. We have that every element of $H$ acts trivially on $\operatorname{{\rm End}}(A)\otimes\mathbb{Z}/2\mathbb{Z}\cong\mathcal{O}_{E}\otimes\mathbb{Z}/2\mathbb{Z}\cong\mathcal{O}_{E}/2\mathcal{O}_{E}\cong\prod_{i}\mathcal{O}_{E}/\mathfrak{p}^{e_{i}}_{i}$. In particular, every element of $H$ acts trivially on each $\mathcal{O}/\mathfrak{p}^{e_{i}}_{i}$, so $H$ is contained in the inertia group $I_{\mathfrak{p}_{i}}$ for every prime $\mathfrak{p}_{i}$ above 2 in $E$. ∎
@@
Let $A/K$ be an abelian variety of dimension $g\geq 1$ with $p=2g+1$ prime. Suppose $\mathfrak{q}$ is a prime of bad reduction for $A$ and there is an element of order $p$ in the image of the inertia group $I_{\mathfrak{q}}$ in $\operatorname{{\rm Gal}}(K(A[\ell^{\prime}])/K)$ for some prime $\ell^{\prime}\neq p$ and $\mathfrak{q}\nmid\ell^{\prime}$. Then either
  • $p$ does not divide $L\colon K]$ and $\operatorname{{\rm End}}^{0}(A)$ is isomorphic to a subfield of $\mathbb{Q}(\zeta_{p})$; or
  • $p$ divides $L\colon K]$, $g\geq 3$ and $A$ is isogenous over $\overline{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield of $\mathbb{Q}(\zeta_{p})$.
@@
Suppose $p$ divides $L\colon\mathbb{Q}$]. Then by [Goo21, Thm. 2.5], $g\geq 3$ and $A$ is isogenous over $\overline{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield $F$ of $\mathbb{Q}(\zeta_{p})$. Hence we may suppose $p$ does not divide $L\colon K]$.
Let $\ell$ be a prime and let $\rho_{\ell}\colon G_{K}\rightarrow\operatorname{{\rm Aut}}_{\mathbb{Z}_{\ell}}(T_{\ell}(A))$ denote the representation of $G_{K}$ arising from its action on the Tate module $T_{\ell}(A)$. By [GR72, Thm. 4.3], for any $\sigma\in I_{\mathfrak{q}}$ and prime $\ell\nmid N(\mathfrak{q})$, the characteristic polynomial of $\rho_{\ell}(\sigma)$ has rational coefficients independent of $\ell$. By assumption, there exists a prime $\ell^{\prime}\neq p$ such that the image of $I_{\mathfrak{q}}$ in $\operatorname{{\rm Gal}}(K(A[\ell^{\prime}])/K)$ contains an element of order $p$. In particular, as $\ell^{\prime}\neq p$, there exists an element $\tau_{\ell^{\prime}}$ in the image of $I_{\mathfrak{q}}$ in $\operatorname{{\rm Aut}}_{\mathbb{Z}_{\ell^{\prime}}}(T_{\ell^{\prime}}(A))$ of order $p$. As the trace of $\tau_{\ell^{\prime}}$ is rational its eigenvalues are the primitive $p$-th roots of unity, owing to the fact $\tau_{\ell^{\prime}}$ has at most 2g=p12g=p-1 distinct eigenvalues. It follows that the characteristic polynomial of $\tau_{\ell^{\prime}}$ is the $p$-th cyclotomic polynomial. Let $\tau\in I_{\mathfrak{q}}$ be such that $\rho_{\ell^{\prime}}(\tau)=\tau_{\ell^{\prime}}$. We deduce that for any prime $\ell\nmid N(\mathfrak{q})$, the characteristic polynomial of $\tau_{\ell}\coloneqq\rho_{\ell}(\tau)$ is the $p$-th cyclotomic polynomial.
By Dirichlet’s theorem on arithmetic progressions, we can find a prime $\ell^{\prime\prime}\nmid N(\mathfrak{q})$, which is a primitive root modulo $p$. The reduction of $\tau_{\ell^{\prime\prime}}$ modulo $\ell^{\prime\prime}$ lands in $\operatorname{{\rm Gal}}(K(A[\ell^{\prime\prime}])/K)$ and has order $p$ thanks to the form of its characteristic polynomial. This allows us to apply [Goo21, Thm. 2.9] and deduce $E\coloneqq\operatorname{{\rm End}}^{0}(A)$ is a field.
As $p$ does not divide $L\colon K]$, for any prime $\ell\nmid N(\mathfrak{q})$, the element $\tau_{\ell}$ lies in the image of $G_{L}$ in $\operatorname{{\rm Aut}}_{\mathbb{Z}_{\ell}}(T_{\ell}(A))$. The decomposition $E\otimes\mathbb{Q}_{\ell}=\prod_{\lambda|\ell}E_{\lambda}$ induces a decomposition $V_{\ell}(A)=\prod_{\lambda|\ell}V_{\lambda}(A)$ giving representations $G_{L}\rightarrow\operatorname{{\rm GL}}_{n}(E_{\lambda})$ where $n=\frac{2g}{[E\colon\mathbb{Q}]}$, see [Rib76, Ch. II] for further details. Let $\tau_{\lambda}\in\operatorname{{\rm GL}}_{n}(E_{\lambda})$ be the projection of $\tau_{\ell}$ onto $V_{\lambda}(A)$. By the above, the eigenvalues of $\tau_{\lambda}$ are distinct primitive $p$-th roots of unity. Taking the trace of $\tau_{\lambda}$ we deduce $E_{\lambda}$ contains a subfield of $\mathbb{Q}(\zeta_{p})$ of degree $E\colon\mathbb{Q}$].
As $\mathbb{Q}(\zeta_{p})$ has a unique subfield $F$ of degree $E\colon\mathbb{Q}$] and for all, but possibly finitely many primes $\lambda$, $E_{\lambda}$ contains $F$, the Chebotarev Density Theorem implies $E$ contains $F$. But $E\colon\mathbb{Q}]=[F\colon\mathbb{Q}$], so in other words, $E\subseteq\mathbb{Q}(\zeta_{p})$. ∎
@@
  • $p$ does not divide $L\colon K]$ and $\operatorname{{\rm End}}^{0}(A)$ is isomorphic to a subfield of $\mathbb{Q}(\zeta_{p})$; or
  • $p$ divides $L\colon K]$, $g\geq 3$ and $A$ is isogenous over $\overline{K}$ to the power of an absolutely simple abelian variety with complex multiplication by a proper subfield of $\mathbb{Q}(\zeta_{p})$.